Graph f(x) = 4x + 1 explain the graph points in words.
Answer by fractalier(2101)

To graph f(x) = 4x + 1 on a coordinate plane, merely choose any arbitrary value for x (usually 2 or 1 or 0 or -1 or -2 work), and see what value the function has for that value of x.
For example, choose x = 1.
The f(1) = 4(1) + 1 = 5.
Thus graph the first point (x, f(x)) as (1, 5).
Then choose x = 2.
Here f(2) = 4(2) + 1 = 9
This point is (2, 9) and so on...